# Timelike conjugate points in Lorentzian length spaces

### v1

@@ 4 Applications
In this section, we give two applications that use of our notions of conjugate points, namely a timelike Rauch comparison theorem (in analogy with Shankar–Sormani [20, Thm. 10.1]), as well as a consequence of the Lorentzian Cartan–Hadamard Theorem established by Erös–Gieger [7]. (The analogous result in the metric setting can be found in Bridson–Haefliger [5]).
@@ 4.1 A timelike Rauch comparison theorem
Let $(X,d,\ll,\leq,\tau)$ be a chronological, regular, locally timelike maximizing and finitely $\tau$-measurable Lorentzian pre-length space with timelike curvature $\leq\kappa$ and let $\gamma\in\operatorname{TGeo}(X)$ be such that $p\coloneqq\gamma(0)$ and $q\coloneqq\gamma(1)$ are symmetric conjugate along $\gamma$. Then $L_{\tau}(\gamma)\geq D_{\kappa}$.
@@
Cover $\gamma([0,1])$ by a finite collection $\mathcal{U}$ of locally timelike maximizing comparison neighborhoods, and let $L$ denote the Lebesgue number of the open cover $\{U\cap\gamma([0,1])\colon U\in\mathcal{U}\}$ of $\gamma([0,1])$. Choose $\delta>0$ such that $d(\gamma(s),\gamma(t))<L$ for any $s,t\in[0,1]$ with $|t-s|<\delta$. Then picking points $0=t_{0}<\dots<t_{N}=1$ in such a way that $|t_{i+1}-t_{i}|<\delta/3$ for each $i$, we can guarantee that each segment $\gamma([t_{i},t_{i+3}])$ lies entirely in one element of $\mathcal{U}$ ($i=0,\dots,N-3$). Because of this, these pieces of $\gamma_{n}$ and $\sigma_{n}$ are maximizing and timelike. Since the endpoints $p,q$ of $\gamma$ are symmetric conjugate, we may find sequences $(\gamma_{n}),(\sigma_{n})\subseteq\operatorname{TGeo}(X)$ with $\gamma_{n}\neq\sigma_{n}$ for all $n$ but with the same endpoints $p_{n},q_{n}$ and $\gamma_{n},\sigma_{n}\to\gamma$ in the $d_{\Gamma}$-topology. It follows that there exist representatives (denoted in the same way) converging to $\gamma$ uniformly and in $\tau$-length. For $n$ large, we therefore also have that, for each $i$, the segments $\gamma_{n}([t_{i},t_{i+3}])$ and $\sigma_{n}([t_{i},t_{i+3}])$ lie in the same element of $\mathcal{U}$ as $\gamma([t_{i},t_{i+3}])$.
In the remainder of the proof we additionally fix $n$ sufficiently large to secure that $\max(L_{\tau}(\gamma_{n}),L_{\tau}(\sigma_{n}))<D_{\kappa}$ and that $p_{n}\ll\gamma_{n}(t_{1})\ll\sigma_{n}(t_{2})\ll\gamma_{n}(t_{3})\ll\dots\ll q_{n}$, as well as $p_{n}\ll\sigma_{n}(t_{1})\ll\gamma_{n}(t_{2})\ll\sigma_{n}(t_{3})\ll\dots\ll q_{n}$.
@@
We may assume that $\bar{\gamma}_{n},\bar{\sigma}_{n}$ are parametrized to unit speed. Consider now a breakpoint $p$ of $\bar{\gamma}_{n}$, and for the moment call $\bar{\alpha}_{n}$ and $\bar{\beta}_{n}$ the geodesic sub-segments of $\bar{\gamma}_{n}$ meeting at $p$. Then replace $\bar{\beta}_{n}$ by its image under an isometry of $\mathbb{L}^{2}(\kappa)$ whose derivative maps the initial velocity vector of $\bar{\beta}_{n}$ to the terminal velocity vector of $\bar{\alpha}_{n}$, thereby straightening the concatenation of $\bar{\alpha}_{n}$ and $\bar{\beta}_{n}$ into an unbroken geodesic segment. We repeat this procedure at each breakpoint of $\bar{\gamma}_{n}$, and analogously for $\bar{\sigma}_{n}$. Concavity of the stacking implies that these unbroken geodesics will remain within the region bounded by the original stacking and therefore also have to meet. Moreover, since we applied an isometry in each step, the resulting unbroken geodesic segments retain the Lorentzian length of $\bar{\gamma}_{n}$ and $\bar{\sigma}_{n}$, respectively. Consequently, we have found two different geodesic segments in $\mathbb{L}^{2}(\kappa)$ that meet at $\bar{\tau}$-length strictly less than $D_{\kappa}$, a contradiction. ∎
@@ 4.2 Nonexistence of conjugate points under nonpositive timelike curvature
Cover $\gamma([0,1])$ by comparison neighborhoods $V_{1},\dots,V_{N}$ such that for a suitable partition $0=t_{0}<\dots<t_{N}=1$, $\gamma([t_{i},t_{i+1}])\subseteq V_{i+1}$ for all $i=0,\dots,N-1$. By (ii), there exists some $k_{0}$ such that also $\gamma_{p_{k}q_{k}}([t_{i},t_{i+1}])\subseteq V_{i+1}$ for all $k\geq k_{0}$ and all $i$. The proof of [7, Prop. 2.3] shows that then each $\gamma_{p_{k}q_{k}}|_{[t_{i},t_{i+1}]}$ as well as $\gamma|_{[t_{i},t_{i+1}]}$ are timelike maximizers. Since $\tau|_{V_{i}\times V_{i}}$ is continuous, we obtain
$L_{\tau}(\gamma_{p_{k}q_{k}}([t_{i},t_{i+1}]))=\tau(\gamma_{p_{k}q_{k}}(t_{i}),\gamma_{p_{k}q_{k}}(t_{i+1}))\to\tau(\gamma(t_{i}),\gamma(t_{i+1}))=L_{\tau}(\gamma|_{[t_{i},t_{i+1}]}).$
@@ 5 Outlook
In this work, we have introduced and studied timelike conjugate points in Lorentzian (pre-)length spaces. We have compared various notions among each other and with their Jacobi field counterpart in the smooth setting. Moreover, we have given applications (Theorems 4.1 and 4.2) illustrating their use.
@@
Appendix A The Fréchet distance of curves
In this Appendix, we introduce the Fréchet distance of curves and prove some basic properties that were used in the main part of the paper. This notion indeed goes back to Fréchet and has found many applications (cf., e.g., [1]), but we are not aware of a reference that rigorously establishes these fundamentals.
Let $(X,d)$ be a metric space. By a path in $X$ we mean an element $\gamma$ of the space of continuous maps $C([0,1],X)$ from $[0,1]$ into $X$ (we use the interval $[0,1]$ for convenience). Two paths $\gamma,\tilde{\gamma}$ are called equivalent if there exists a parametrization $\varphi\colon[0,1]\to[0,1]$ which is continuous, strictly increasing, and surjective such that $\gamma=\tilde{\gamma}\circ\varphi$. A curve is an equivalence class of paths under this equivalence relation, and we will write $\gamma]$ for the curve represented by the path $\gamma$. A path is called nowhere constant if it is not constant on any non-trivial sub-interval of $[0,1]$. This notion clearly descends to the space of curves as well. Given two curves $\gamma],[\sigma]$, we define the Fréchet distance between them as follows:
@@ Appendix A The Fréchet distance of curves
Let $\alpha\colon[0,1]\to[0,1]^{2}$ be a continuous, coordinate-wise non-decreasing map given by $\alpha(t)=(\varphi(t),\psi(t))$ and let $\alpha(0)=(0,0)$ and $\alpha(1)=(1,1)$. Then there exists a coordinate-wise non-decreasing and $1$-Lipschitz continuous curve $\tilde{\alpha}=(\tilde{\varphi},\tilde{\psi})\colon[0,2]\to[0,1]^{2}$ that has the same image as $\alpha$.
@@
Finally, to show that $\tilde{\alpha}$ is $1$-Lipschitz, let $0\leq u<v\leq 2$. Then using monotonicity of $\theta$, $\varphi$ and $\psi$, hence of $\tilde{\varphi}$ and $\tilde{\psi}$,
@@
Let $N$ be the set of continuous, componentwise non-decreasing maps $\alpha=(\varphi,\psi)\colon[0,1]\to[0,1]^{2}$ with $\alpha(0)=(0,0)$ and $\alpha(1,1)=(1,1)$. Then $\mathcal{N}\coloneqq\{\alpha([0,1])\mid\alpha\in N\}$ is compact with respect to the Hausdorff distance $d_{H}$.
@@
Denote by $L$ the set of all coordinate-wise non-decreasing $1$-Lipschitz maps $\eta\colon[0,2]\to[0,1]^{2}$ with $\eta(0)=(0,0)$ and $\eta(1)=(1,1)$. By the Theorem of Arzelà-Ascoli, $L$ is relatively compact in $C([0,2],[0,1]^{2})$. Since it is clearly also closed with respect to uniform convergence, it is indeed compact in $C([0,2],[0,1]^{2})$.
@@
Since $d_{H}(\iota(\eta),\iota(\zeta))\leq\sup_{u\in[0,2]}\|\eta(u)-\zeta(u)\|_{2}$, it follows that $\iota$ is continuous (indeed, $1$-Lipschitz). Thus $\mathcal{L}=\iota(L)$ is compact with respect to $d_{H}$. Finally, by Lemma A.1, $\mathcal{N}=\mathcal{L}$, concluding the proof. ∎
@@ Acknowledgments
The authors would like to thank Christina Sormani for helpful discussions.
This research was funded in part by the Austrian Science Fund (FWF) [Grant DOI 10.55776/EFP6 and 10.55776/J4913]. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.

### v2

@@ 4 Applications
In this section, we give two applications that use of our notions of conjugate points, namely a timelike Rauch comparison theorem (in analogy with Shankar–Sormani [21, Thm. 10.1]), as well as a consequence of the Lorentzian Cartan–Hadamard Theorem established by Erös–Gieger [7] (for the metric Cartan–Hadamard Theorem, cf. Bridson–Haefliger [5, Thm. 4.1]).
@@ 4.1 A timelike Rauch comparison theorem
Let $(X,d,\ll,\leq,\tau)$ be a chronological, regular and locally timelike maximizing Lorentzian pre-length space with timelike curvature $\leq\kappa$ and let $\gamma\in\operatorname{TGeo}(X)$ be such that $p\coloneqq\gamma(0)$ and $q\coloneqq\gamma(1)$ are symmetric conjugate along $\gamma$. Then $L_{\tau}(\gamma)\geq D_{\kappa}$.
@@
Cover $\gamma([0,1])$ by a finite collection $\mathcal{U}$ of locally timelike maximizing comparison neighborhoods, and let $L$ denote the Lebesgue number of the open cover $\{U\cap\gamma([0,1])\colon U\in\mathcal{U}\}$ of $\gamma([0,1])$. Choose $\delta>0$ such that $d(\gamma(s),\gamma(t))<L$ for any $s,t\in[0,1]$ with $|t-s|<\delta$. Then picking points $0=t_{0}<\dots<t_{N}=1$ in such a way that $|t_{i+1}-t_{i}|<\delta/3$ for each $i$, we can guarantee that each segment $\gamma([t_{i},t_{i+3}])$ lies entirely in one element of $\mathcal{U}$ ($i=0,\dots,N-3$).
Figure 1: Geometric setup in the proof of Theorem 4.1
Since the endpoints $p,q$ of $\gamma$ are symmetric conjugate, we may find sequences $(\gamma_{n}),(\sigma_{n})\subseteq\operatorname{TGeo}(X)$ with $\gamma_{n}\neq\sigma_{n}$ for all $n$ but with the same endpoints $p_{n},q_{n}$ and $\gamma_{n},\sigma_{n}\to\gamma$ in the $d_{\Gamma}$-topology. It follows that there exist representatives (denoted in the same way) converging to $\gamma$ uniformly and in $\tau$-length. For $n$ large, we therefore also have that, for each $i$, the segments $\gamma_{n}([t_{i},t_{i+3}])$ and $\sigma_{n}([t_{i},t_{i+3}])$ lie in the same element of $\mathcal{U}$ as $\gamma([t_{i},t_{i+3}])$.
In the remainder of the proof we additionally fix $n$ sufficiently large to secure that $\max(L_{\tau}(\gamma_{n}),L_{\tau}(\sigma_{n}))<D_{\kappa}$ and that $p_{n}\ll\gamma_{n}(t_{1})\ll\sigma_{n}(t_{2})\ll\gamma_{n}(t_{3})\ll\dots\ll q_{n}$, as well as $p_{n}\ll\sigma_{n}(t_{1})\ll\gamma_{n}(t_{2})\ll\sigma_{n}(t_{3})\ll\dots\ll q_{n}$ (cf. Figure 1).
@@
We may assume that $\bar{\gamma}_{n},\bar{\sigma}_{n}$ are parametrized to unit speed. Consider now a breakpoint $p$ of $\bar{\gamma}_{n}$, and for the moment call $\bar{\alpha}_{n}$ and $\bar{\beta}_{n}$ the geodesic sub-segments of $\bar{\gamma}_{n}$ meeting at pp. Then replace $\bar{\beta}_{n}$ by its image under an isometry of $\mathbb{L}^{2}(\kappa)$ whose derivative maps the initial velocity vector of $\bar{\beta}_{n}$ to the terminal velocity vector of $\bar{\alpha}_{n}$, thereby straightening the concatenation of $\bar{\alpha}_{n}$ and $\bar{\beta}_{n}$ into an unbroken geodesic segment. We repeat this procedure at each breakpoint of $\bar{\gamma}_{n}$, and analogously for $\bar{\sigma}_{n}$. Concavity of the stacking implies that these unbroken geodesics will remain within the region bounded by the original stacking and therefore also have to meet. Moreover, since we applied an isometry in each step, the resulting unbroken geodesic segments retain the Lorentzian length of $\bar{\gamma}_{n}$ and $\bar{\sigma}_{n}$, respectively. Consequently, we have found two different geodesic segments in $\mathbb{L}^{2}(\kappa)$ that meet at $\bar{\tau}$-length strictly less than $D_{\kappa}$, a contradiction. ∎
@@ 4.2 Nonexistence of conjugate points under nonpositive timelike curvature
Cover $\gamma([0,1])$ by comparison neighborhoods $V_{1},\dots,V_{N}$ such that for a suitable partition $0=t_{0}<\dots<t_{N}=1$, $\gamma([t_{i},t_{i+1}])\subseteq V_{i+1}$ for all $i=0,\dots,N-1$. By (ii), there exists some $k_{0}$ such that also $\gamma_{p_{k}q_{k}}([t_{i},t_{i+1}])\subseteq V_{i+1}$ for all $k\geq k_{0}$ and all ii. The proof of [7, Prop. 2.3] shows that then each $\gamma_{p_{k}q_{k}}|_{[t_{i},t_{i+1}]}$ as well as $\gamma|_{[t_{i},t_{i+1}]}$ are timelike maximizers. Since $\tau|_{V_{i}\times V_{i}}$ is continuous, we obtain
$L_{\tau}(\gamma_{p_{k}q_{k}}|_{[t_{i},t_{i+1}]})=\tau(\gamma_{p_{k}q_{k}}(t_{i}),\gamma_{p_{k}q_{k}}(t_{i+1}))\to\tau(\gamma(t_{i}),\gamma(t_{i+1}))=L_{\tau}(\gamma|_{[t_{i},t_{i+1}]}).$
@@ 5 Outlook
In this work, we have introduced and studied timelike conjugate points in Lorentzian (pre-)length spaces. We have compared various notions with one another and with their Jacobi field counterpart in the smooth setting. Moreover, we have given applications (Theorems 4.1 and 4.2) illustrating their use.
@@
Appendix A The Fréchet distance between curves
In this Appendix, we introduce the Fréchet distance between curves and prove some basic properties that were used in the main part of the paper. This notion indeed goes back to Fréchet and has found many applications (cf., e.g., [1]), but we are not aware of a reference that rigorously establishes these fundamentals.
Let $(X,d)$ be a metric space. By a path in $X$ we mean an element $\gamma$ of the space of continuous maps $C([0,1],X)$ from $[0,1]$ into $X$ (we use the interval $[0,1]$ for convenience). Two paths $\gamma,\tilde{\gamma}$ are called equivalent if there exists a parametrization $\varphi\colon[0,1]\to[0,1]$ which is strictly increasing and surjective (hence continuous) such that $\gamma=\tilde{\gamma}\circ\varphi$. A curve is an equivalence class of paths under this equivalence relation, and we will write $\gamma]$ for the curve represented by the path $\gamma$. A path is called nowhere constant if it is not constant on any non-trivial sub-interval of $[0,1]$. This notion clearly descends to the space of curves as well. Given two curves $\gamma],[\sigma]$, we define the Fréchet distance between them as follows:
@@ Appendix A The Fréchet distance between curves
Let $\alpha\colon[0,1]\to[0,1]^{2}$ be a continuous, coordinate-wise non-decreasing map given by $\alpha(t)=(\varphi(t),\psi(t))$ with $\alpha(0)=(0,0)$ and $\alpha(1)=(1,1)$. Then there exists a coordinate-wise non-decreasing and 11-Lipschitz continuous curve $\tilde{\alpha}=(\tilde{\varphi},\tilde{\psi})\colon[0,2]\to[0,1]^{2}$ that has the same image as $\alpha$.
@@
Finally, to show that $\tilde{\alpha}$ is 11-Lipschitz, let $0\leq u<v\leq 2$. Then using monotonicity of $\theta$, $\varphi$ and $\psi$, hence of $\tilde{\varphi}$ and $\tilde{\psi}$,
@@
Let NN be the set of continuous, coordinate-wise non-decreasing maps $\alpha=(\varphi,\psi)\colon[0,1]\to[0,1]^{2}$ with $\alpha(0)=(0,0)$ and $\alpha(1)=(1,1)$. Then $\mathcal{N}\coloneqq\{\alpha([0,1])\mid\alpha\in N\}$ is compact with respect to the Hausdorff distance $d_{H}$.
@@
Denote by $L$ the set of all coordinate-wise non-decreasing 11-Lipschitz maps $\eta\colon[0,2]\to[0,1]^{2}$ with $\eta(0)=(0,0)$ and $\eta(1)=(1,1)$. By the Theorem of Arzelà–Ascoli, $L$ is relatively compact in $C([0,2],[0,1]^{2})$. Since it is clearly also closed with respect to uniform convergence, it is indeed compact in $C([0,2],[0,1]^{2})$.
@@
Since $d_{H}(\iota(\eta),\iota(\zeta))\leq\sup_{u\in[0,2]}\|\eta(u)-\zeta(u)\|_{2}$, it follows that $\iota$ is continuous (indeed, 11-Lipschitz). Thus $\mathcal{L}=\iota(L)$ is compact with respect to $d_{H}$. Finally, by Lemma A.1, $\mathcal{N}=\mathcal{L}$, concluding the proof. ∎
@@ Acknowledgments
The authors would like to thank Christina Sormani and Nathalie Tassotti for helpful discussions.
This research was funded in part by the Austrian Science Fund (FWF) [Grant DOI 10.55776/EFP6 and 10.55776/J4913]. For open access purposes, the authors have applied a $C$ BY public copyright license to any author accepted manuscript version arising from this submission.